OpenStudy (anonymous):

What is the relationship between the lines determined by the following two equations? 6x – 2y = 16 y = 3x – 8 A. they are the same line B. parallel C. neither parallel nor perpendicular D. perpendicular

OpenStudy (shamil98):

6x - 2y = 16 y = 3x - 8 -2y = -6x + 16 y = 3x - 8 so they are the same line.